40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jan 23 15:16:52 2017, gautam, Update, Optical Levers, ETMY Oplev HeNe needs to be replaced ETMY_Oplev.pdf
    Reply  Mon Jan 23 17:24:26 2017, Steve, Update, Optical Levers, ETMY Oplev HeNe is running hot ETMY_oplev__lasers.png
       Reply  Tue Jan 24 17:52:15 2017, Craig, Update, Optical Levers, ETMY Oplev HeNe is replaced 6x
          Reply  Wed Mar 1 08:28:04 2017, Steve, Update, Optical Levers, oplev laser RIN RIN_1103P.pngRIN_HeNe.png
             Reply  Thu Mar 2 13:59:04 2017, Steve, Update, Optical Levers, oplev laser RIN 1103P@2d.png
                Reply  Thu Mar 2 17:58:45 2017, rana, Update, Optical Levers, oplev laser RIN 
                   Reply  Fri Mar 3 17:24:21 2017, Steve, Update, Optical Levers, oplev laser RIN 3march17.pdf
                      Reply  Wed Mar 8 15:28:37 2017, Steve, Update, Optical Levers, oplev laser RIN RIN.jpgRIN_1103P_rotated.pdf
                      Reply  Mon Apr 10 09:49:32 2017, Steve, Update, Optical Levers, oplev laser RIN test planning 
                         Reply  Mon Apr 10 14:21:57 2017, rana, Update, Optical Levers, oplev laser RIN test planning 
       Reply  Tue May 30 13:36:58 2017, Steve, Update, Optical Levers, ETMY Oplev HeNe is replaced oplevETMY120d.png
          Reply  Tue Jan 15 18:16:00 2019, gautam, Update, Optical Levers, ETMY Oplev HeNe needs replacement OLRIN.pdfOLsums.png
             Reply  Fri Jan 18 12:52:07 2019, gautam, Update, Optical Levers, BS/PRM Oplev HeNe replaced OLRIN_20190118.pdf
Message ID: 14402     Entry time: Tue Jan 15 18:16:00 2019     In reply to: 13018     Reply to this: 14404
Author: gautam 
Type: Update 
Category: Optical Levers 
Subject: ETMY Oplev HeNe needs replacement 

Perhaps the ETMY Oplev HeNe is also giving up - the power has fallen by ~30% over 1 year (Attachment #2), nearly twice as much as ETMX but the RIN spectrum (Attachment #1, didn't even need to rotate it!) certainly seems suspicious. Some "nominal" RIN levels for HeNes can be found earlier in this thread. I can't close any of the EY Oplev loops in this condition. I'll double check to make sure I'm routing the right beam onto the QPD, but if the problem persists, I'll replace the HeNe. ITMX HeNe also looks to be near EOL.

Quote:

Finally I reallized what is killing the ETMY oplev laser. Wrong  power supply, it  was driving the HeNe laser by 600V higher voltage than recommended. Power supply 101T-2300Vdc replaced by 101T-1700Vdc ( Uniphase model 1201-1, sn 2712420 )

The laser head 1103P, sn P947049 lived for 120 days and it was replaced by sn P964431   New laser output 2.8 mW,  quadrant sum 19,750 counts

Attachment 1: OLRIN.pdf  68 kB  | Hide | Hide all
OLRIN.pdf
Attachment 2: OLsums.png  44 kB  | Hide | Hide all
OLsums.png
ELOG V3.1.3-